OpenStudy (anonymous):

is the derivative of y=3 squareroot of (2t over (t squared - 4)) ... (2t to the onethird over 2t) ??? please help

OpenStudy (anonymous):

\[the problem is y= \sqrt[3]{ \frac{ 2t }{ t ^{2} -4}}\]

OpenStudy (anonymous):

you can also write it in the form: \[y=\left( \frac{ 2t }{ t ^{2}-4 } \right)^{\frac{ 1 }{ 3 }}\] then you can easily apply the chain rule.

OpenStudy (anonymous):

can you please give me a minute.. im going to try that and can you let me know if i'm right please

OpenStudy (anonymous):

sure:) np.

OpenStudy (anonymous):

can you let me know if im going in the right direction .. sorry im like slow >,< \[\frac{ 1 }{ 3 }\left(\begin{matrix}\frac{ 2t }{ t ^{2}-4 } \\\end{matrix}\right)^{\frac{ -2 }{ 3 }} \times \left(\begin{matrix} \\ \frac{ 2 }{ 2t }\end{matrix}\right)\]

OpenStudy (anonymous):

o wait no.. i made things more complicated

OpenStudy (anonymous):

the first part is right, but the derivative of the inside function is not right, because you need to use the quotient rule in finding the derivative.. no problem, i'm slow learner too :)

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

lol thanks

OpenStudy (anonymous):

h.o

OpenStudy (anonymous):

you got it?

OpenStudy (anonymous):

i hope this is right. \[-2t ^{2}-8 ?\]

OpenStudy (anonymous):

i did \[(t ^{2}-4)(2) - (2t)(2t)\]

OpenStudy (anonymous):

for the derivative of the inside i mean

OpenStudy (anonymous):